# GenAnalysis: Joint Shape Analysis by Learning Man-Made Shape Generators with Deformation Regularizations

### v1

@@ 1. Introduction
Shape analysis, is a fundamental research area in geometry processing that enjoys many applications (Mitra et al., 2014; Xu et al., 2017). Existing shape analysis approaches fall into the categories of supervised and unsupervised methods. Although supervised methods achieved state-of-the-art results, they require user labels, which are costly to obtain and do not scale up. In contrast, unsupervised shape analysis methods (Deng et al., 2021; Shuai et al., 2023; Chen et al., 2024) benefit from learning from large-scale unlabeled shapes and have recently shown promising results in the large-scale ShapeNetPart (Yi et al., 2016) dataset.
The fundamental challenge in developing unsupervised techniques is to accommodate significant structural and geometric variations across the input shapes. A popular paradigm is learning template-based models in which each template consists of a bag of (non-rigidly) deforming parts. The learning phase learns parameters of the templates to fit the training data. These include early works of (Ovsjanikov et al., 2011; Kim et al., 2013; Averkiou et al., 2014; Huang et al., 2015a, b) and recent deep learning methods (Ganapathi-Subramanian et al., 2018; Genova et al., 2019a; Deprelle et al., 2019; Shuai et al., 2023; Chen et al., 2024; Zheng et al., 2021; Deng et al., 2021). The key advantage of template-based models is that during the testing phase it is only required to fit template to each shape, which is efficient and simple. However, template-based models exhibit two major limitations. First, template-based models rely on prior knowledge of shape parts (e.g., the number of parts and their spatial locations) either explicitly or implicitly. Second, the learning and inference phases can easily get stuck in local minimums.
In this paper, we introduce a novel shape analysis framework, named GenAnalysis, which learns an implicit shape generative model to fit the training shapes. We will show how to use this template-free shape generative model to perform shape analysis, hence circumventing the limitation of template-based models. Our approach is inspired by the success of studying organic shape collections using the lens of shape spaces as differential manifolds (Klassen et al., 2004; Kilian et al., 2007; Srivastava et al., 2011; Hartman et al., 2023; Yang et al., 2023a). Moreover, recent advances in shape generative models (Park et al., 2019; Chen and Zhang, 2019; Mescheder et al., 2019; Atzmon and Lipman, 2021; Yang et al., 2024) can accurately fit man-made training shapes that exhibit much larger geometric variations than organic shapes. These learned continuous man-made shape spaces offer novel means of shape analysis. Specifically, they provide smooth interpolations between two shapes that can be used to compute shape correspondences. Moreover, when trained properly,these generative models enable us to understand the shape variations among the shape collection for shape segmentation by analyzing shape variations in the tangent space of each shape when viewing the learned generative models as shape manifolds. This continuous shape-space approach greatly alleviates the local minimum issue in template-based approaches.
@@
The proposed loss builds on the popular piece-wise affine assumption (Xu et al., 2010; Ovsjanikov et al., 2011; Kim et al., 2013; Huang et al., 2015a; Chen et al., 2024), in which if we abstract each shape part using a bounding box, then the deformations at the part level are affine (see Figure 2). However, the challenge is how to model the significant shape variations within each part. GenAnalysis presents two strategies to address this challenge. First, we minimize the piece-wise affine deformation between adjacent synthetic shapes defined by the generator (i.e., they are close in the latent space). This approach is analogous to the smoothness loss in curve fitting, but we model strong prior knowledge about inter-shape deformations. In particular, we show how to employ an as-affine-as-possible deformation loss under a robust norm to fulfill the piece-wise affine assumption. Second, we describe a weighting scheme for test-time optimization (Wang et al., 2021; Chen et al., 2022; Niu et al., 2022b; Wang et al., 2023). It prioritizes that the regularization loss distributes deviations of piece-wise affine deformations away from shapes of interest. This allows us to extract segmentation cues by analyzing the tangent space of each shape (i.e., vector fields that characterize differences between each shape and its adjacent shapes in the latent space).
@@
Figure 2. Piece-wise affine assumption. Each shape part from source shape, approximated by its bounding box, undergoes a separate affine transformation $A_{i}$ to the corresponding part in the target shape.
@@ 4.1.1. AAAP deformation induced correspondences
To this end, we first discretize $g^{\theta}(\boldsymbol{x},\boldsymbol{z})=0$ using a mesh with $n\approx 2000$ vertices $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}),1\leq i\leq n$, e.g., via Marching Cube (Lorensen and Cline, 1987). To define the AAAP deformation model, our goal is to determine the corresponding location $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v})$ on the neighboring implicit surface $g^{\theta}(\boldsymbol{x},\boldsymbol{z}+\epsilon\boldsymbol{v})=0$, where $\boldsymbol{v}$ is the direction of the perturbation and $\epsilon=10^{-3}$ is an infinitesimal value. We approximate $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v})$ via the first-order Taylor expansion $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v})\approx\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z})+\epsilon\boldsymbol{d}_{i}^{\boldsymbol{v}}(\boldsymbol{z})$.By computing the derivative of the implicit function $g^{\theta}(\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v}),\boldsymbol{z}+\epsilon\boldsymbol{v})=0$ with respect to $\boldsymbol{v}$ using the chain rule, we arrive at the following linear constraint of $\boldsymbol{d}_{i}^{\boldsymbol{v}}(\boldsymbol{z})$:
@@ 4.1.3. Test-Time Optimization
Figure 8. Interpolations between the left shape and the right shape that using a path of two intermediate shapes. The underlying deformation deviates from the piece-wise affine assumption. We minimize AAAP deformations between adjacent shapes. (Top) Using uniform weights $(1,1,1)$ for three shape pairs. The deformations of all shape pairs deviate from piece-wise affine. (Bottom) The weights are $(1,0.1,0.1)$ where the first pair has a large weight. The deformation of the first pair is close to piece-wise affine.
@@ 4.3. Shape Matching
Figure 10. Correspondences computation of samples from the source shape through intermediate shapes to the target shape. Corresponding points share the same colors. We alternate between a propagation step and a projection step. Without projection, the correspondences do not lie on the target shape and also drift away.
@@
Figure 11. Color-map of similarity weights $w(\boldsymbol{p}_{i})$ between the source shape (left) and the target shape (right) defined by the local distortions $e_{i}$ of the correspondences computed using our approach. We can see that the similarity weights characterize the structural similarities.
@@ 4.4. Consistent Segmentation
Figure 13. Comparison between over-segments derived from different strategies. (Top) Over-segments derived from angles between adjacent faces (Golovinskiy and Funkhouser, 2008). (Bottom) Over-segments using our approach, which offer more meaningful segment boundaries.
@@ 5.2.2. Keypoint transfer
Table 3. Shape co-segmentation result on ShapeNetPart. We report the mean IOU. Higer is better. Baseline approaches include BAE-Net (Chen et al., 2019),RIM-Net (Niu et al., 2022a), and DAE-Net (Chen et al., 2024).
@@ 5.4.1. Without regularization
GenAnalysis-No-AAAPReg 58.1 49.2 52.4
GenAnalysis-No-TanAnal. 72.3 69.2 62.4
GenAnalysis-No-TestTime. 85.7 79.2 74.3
GenAnalysis-No-Weighting. 87.6 81.5 77.6
@@
Table 4. Shape co-segmentation result on ShapeNetPart. We report the mean IOU. Higer is better. GenAnalysis-AAAPReg means GenAnalysis without AAAP regularization. GenAnalysis-No-TanAnal means no segmentation cues from tangent space analysis. GenAnalysis-No-Testime stands for dropping test-time optimization. GenAnalysis-No-Weighting means dropping the latent code weighting term in test time optimization.
@@ Appendix A Expression of the Quadratic Energy
Here $K=\mathrm{diag}(\{|\mathcal{N}_{i}|\})\otimes I_{3}$; $E^{\theta}(\boldsymbol{z})$ and $G^{\theta}(\boldsymbol{z})$ are $n\times n$ sparse and diagonal block matrices, whose expressions are deferred to Appendix A.1.
@@ A.3. Correspondence Computation
Appendix B Spectral Consistent Segmentation
$\overline{L}=I_{nm}-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}.$
We form an embedding space
Appendix C Consistent Segmentation Algorithm Detail
We summarize the algorithm presented in section 4.4 in algorithm block.
@@ Appendix C Consistent Segmentation Algorithm Detail
Appendix D More qualitative results
We show additional consistent segmentation results on ShapeNet in Figure  20,  21,  22,  23 respectively. We show more results on part label transfer and key point transfer in Figure  24 and  25.
@@ Appendix D More qualitative results
Figure 20. More qualitative results of shape co-segmentation on chair category in ShapeNet.
@@
Figure 21. More qualitative results of shape co-segmentation on table category in ShapeNet.
@@
Figure 22. More qualitative results of shape co-segmentation on airplane category in ShapeNet.
[…]
Figure 23. More qualitative results of shape co-segmentation on other category in ShapeNet.
@@
Figure 25. Additional comparison on keypoint transfer in ShapeNet.

### v2

@@ 1. Introduction
Shape analysis is a fundamental research area in geometry processing that enjoys many applications (Mitra et al., 2014; Xu et al., 2017). Existing shape analysis approaches fall into the categories of supervised and unsupervised methods. Although supervised methods achieved state-of-the-art results, they require user labels, which are costly to obtain and do not scale up. In contrast, unsupervised shape analysis methods (Deng et al., 2021; Shuai et al., 2023; Chen et al., 2024) benefit from learning from large-scale unlabeled shapes and have recently shown promising results in the large-scale ShapeNetPart (Yi et al., 2016) dataset.
@@
The fundamental challenge in developing unsupervised techniques is to accommodate significant structural and geometric variations across the input shapes. A popular paradigm is learning template-based models in which each template consists of a bag of (non-rigidly) deforming parts. The learning phase learns parameters of the templates to fit the training data. These include early works of (Ovsjanikov et al., 2011; Kim et al., 2013; Averkiou et al., 2014; Huang et al., 2015a, b) and recent deep learning methods (Ganapathi-Subramanian et al., 2018; Genova et al., 2019a; Deprelle et al., 2019; Shuai et al., 2023; Chen et al., 2024; Zheng et al., 2021; Deng et al., 2021). The key advantage of template-based models is that during the testing phase it is only required to fit template to each shape, which is efficient and simple. However, template-based models exhibit two major limitations. First, template-based models rely on strong prior knowledge of shape parts (e.g., the number of parts and their spatial locations) either explicitly or implicitly. Second, the learning and inference phases can easily get stuck in local minimums.
In this paper, we introduce a novel shape analysis framework, named GenAnalysis, which learns an implicit shape generative model to fit the training shapes. We will show how to use this template-free shape generative model to perform shape analysis, hence circumventing the limitation of template-based models. Our approach is inspired by the success of studying organic shape collections using the lens of shape spaces as differential manifolds (Klassen et al., 2004; Kilian et al., 2007; Srivastava et al., 2011; Hartman et al., 2023; Yang et al., 2023a). Moreover, recent advances in shape generative models (Park et al., 2019; Chen and Zhang, 2019; Mescheder et al., 2019; Atzmon and Lipman, 2021; Yang et al., 2024) can accurately fit man-made training shapes that exhibit much larger geometric variations than organic shapes. These learned continuous man-made shape spaces offer novel means of shape analysis. Specifically, they provide smooth interpolations between two shapes that can be used to compute shape correspondences. Moreover, when trained properly,these generative models enable us to understand the shape variations among the shape collection for shape segmentation by analyzing shape variations in the tangent space of each shape when viewing the learned generative models as shape manifolds.
@@
The proposed loss builds on the popular piece-wise affine assumption (Xu et al., 2010; Ovsjanikov et al., 2011; Kim et al., 2013; Huang et al., 2015a; Chen et al., 2024), in which if we abstract each shape part using a bounding box, then the deformations at the part level are affine (see Figure 2). It also enforces local piece-wise affine assumption in a more flexible and faithful way compared to template based approaches hence alleviates the local minimum issue by achieving better shape quality. However, the challenge is how to model the significant shape variations within each part. GenAnalysis presents two strategies to address this challenge. First, we minimize the piece-wise affine deformation between adjacent synthesized shapes defined by the generator (i.e., they are close in the latent space). This approach is analogous to the smoothness loss in curve fitting, but we model strong prior knowledge about inter-shape deformations. In particular, we show how to employ an as-affine-as-possible deformation loss under a robust norm to fulfill the piece-wise affine assumption. Second, we describe a weighting scheme for test-time optimization (Wang et al., 2021; Chen et al., 2022; Niu et al., 2022b; Wang et al., 2023). It prioritizes that the regularization loss distributes deviations of piece-wise affine deformations away from shapes of interest. This allows us to extract segmentation cues by analyzing the tangent space of each shape (i.e., vector fields that characterize differences between each shape and its adjacent shapes in the latent space).
@@
Figure 2. Piece-wise affine assumption. Each shape part from the source shape, approximated by its bounding box, undergoes an affine transformation $A_{i}$ to the corresponding part in the target shape.
[…]
Two chair silhouettes divided into parts; arrows show each part mapped by its own affine transform $A_{i}$.
@@ 4.1.1. AAAP deformation induced correspondences
To this end, we first discretize $g^{\theta}(\boldsymbol{x},\boldsymbol{z})=0$ using a mesh with $n\approx 2000$ vertices $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}),1\leq i\leq n$, e.g., via Marching Cube (Lorensen and Cline, 1987). To define the AAAP deformation model, our goal is to determine the corresponding location $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v})$ on the neighboring implicit surface $g^{\theta}(\boldsymbol{x},\boldsymbol{z}+\epsilon\boldsymbol{v})=0$, where $\boldsymbol{v}$ is the direction of the perturbation and $\epsilon=10^{-3}$ is an infinitesimal value. We approximate $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v})$ via the first-order Taylor expansion $\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v})\approx\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z})+\epsilon\boldsymbol{d}_{i}^{\boldsymbol{v}}(\boldsymbol{z})$.
By computing the derivative of the implicit function $g^{\theta}(\boldsymbol{p}_{i}^{\theta}(\boldsymbol{z}+\epsilon\boldsymbol{v}),\boldsymbol{z}+\epsilon\boldsymbol{v})=0$ with respect to $\boldsymbol{v}$ using the chain rule, we arrive at the following linear constraint of $\boldsymbol{d}_{i}^{\boldsymbol{v}}(\boldsymbol{z})$:
@@ 4.1.3. Test-Time Optimization
Figure 8. Interpolations between the left and right shapes via two intermediates. We minimize AAAP deformations between adjacent shapes. (Top) Uniform weights $(1,1,1)$ cause all deformations to deviate from piece-wise affine. (Bottom) Weights $(1,0.1,0.1)$ emphasize the first pair, yielding a near–piecewise‐affine first deformation.
[…]
Six small meshes arranged in two rows of three. The top row shows uniform–weight deformations deviating from piece‐wise affine; the bottom row shows first‐pair‐weighted deformations that closely match the affine assumption.
@@ 4.3. Shape Matching
Figure 10. Correspondence computation from the source shape through intermediate shapes to the target. Points share the same colours as they move; alternating propagation and projection steps keep them on the surface. Without projection, correspondences drift away and fail to lie on the target.
[…]
A four-panel diagram showing source shape, propagation step, projection step, and target shape, with colour‐matched points and labels.
@@
Figure 11. Color‑map of similarity weights $w(\boldsymbol{p}_{i})$ between the source shape (left) and the target shape (right), defined by the local distortions $e_{i}$ of the correspondences computed using our approach. The similarity weights clearly characterize structural similarity across the shapes.
[…]
Two silhouettes side‑by‑side colored according to similarity weights: cooler colours indicate low distortion, warmer colours indicate high distortion.
@@ 4.4. Consistent Segmentation
Figure 13. Comparison of over‑segmentations from different methods. (Top) Segments based on angles between adjacent faces (Golovinskiy and Funkhouser, 2008). (Bottom) Segments produced by our approach, which yield more meaningful boundaries.
[…]
Two rows of segmented shapes: the top row shows angular‐based over‑segmentation, the bottom row shows our improved segment boundaries.
@@ 5.2.2. Keypoint transfer
Table 3. Shape co-segmentation results on ShapeNetPart. We report mean IoU (higher is better). Baselines include BAE‑Net (Chen et al., 2019), RIM‑Net (Niu et al., 2022a), and DAE‑Net (Chen et al., 2024).
@@ 5.4.1. Without regularization
Table 4. Shape co-segmentation results on ShapeNetPart. We report the mean IoU (higher is better). “GenAnalysis–NR” omits AAAP regularization; “GenAnalysis–No–TanAnal” drops tangent‑space segmentation cues; “GenAnalysis–No–TestTime” removes test‑time optimization; “GenAnalysis–No–Weighting” omits latent‑code weighting at test time.
@@
GenAnalysis–NR 58.1 49.2 52.4
GenAnalysis–No–TanAnal 72.3 69.2 62.4
GenAnalysis–No–TestTime 85.7 79.2 74.3
GenAnalysis–No–Weighting 87.6 81.5 77.6
@@ 6. Conclusions, Limitations, and Future Work
Acknowledgements.
Yuezhi Yang and Qixing Huang were supported by NSF IIS-2047677, NSF IIS-2413161, and gifts from Adobe and Google.
@@ Appendix A Expression of the Quadratic Energy
Here $K=\mathrm{diag}(\{|\mathcal{N}_{i}|\})\otimes I_{3}$; $E^{\theta}(\boldsymbol{z})$ and $G^{\theta}(\boldsymbol{z})$ are $n\times n$ sparse diagonal block matrices, whose expressions are deferred to Appendix A.1.
@@ A.3. Correspondence Computation
Appendix B Spectral Consistent Segmentation
$\overline{L}=I_{nm}-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}.$
We form an embedding space
Appendix C Consistent Segmentation Algorithm Detail
We summarize the consistent segmentation algorithm presented in section 4.4 in algorithm 1.
Table 5. Reconstruction error on three categories of ShapeNetPart. We report chamfer distance at scale 1e-3. Lower is better.
CD (in 1e-3 scale) Chair Table Airplane
BAE (Deprelle et al., 2019) 1.316 1.915 0.593
RIM (Genova et al., 2019b) 1.320 1.699 0.630
DAE (Zheng et al., 2021) 1.244 1.579 0.328
Ours 0.865 1.246 0.235
Appendix D Reconstruction Quality Evaluation
We evaluate the representation capability of learned latent space of GenAnalysis by measuring the reconstruction error using the chamfer distance (CD) in Table  5. Figure 14 shows qualitative comparison of the reconstruction quality against baseline methods. Our results indicate a clear improvement over the baseline methods, since our approach avoids model expressivity limitations commonly associated with template-fitting techniques.
Appendix E Texture Transfer Comparison
We compare texture transfer result using our method and baseline methods including DIF(Deng et al., 2021), DIT(Zheng et al., 2021) and semantic DIF(Kim et al., 2023). A qualitative comparison is given in Figure 20. Our results show better preservation in fine details and overall patterns of transfered texture compared to baseline approaches.
@@ Appendix E Texture Transfer Comparison
Figure 20. Texture transfer result in ShapeNet.
Appendix F Segmentation Granularity Ablation
We present segmentation result at different level of granularity in Figure  21 by using different number of cluster K in spectral consistent segmentation algorithm, as detailed in appendix B. We achieve finer level segmentation by increasing number of clusters. In all other experiment, we adjust hyper-parameters of clustering so that the resulting segments are at the same level of granularity of baseline approaches.
Appendix G More qualitative results
We show additional result on consistent segmentation, part label transfer and key point transfer in Figure  24,  23 and  22.
@@ Appendix G More qualitative results
Figure 21. Ablation study on number of cluster K used in spectral consistent segmentation algorithm on earphone and guitar category.
@@
Figure 22. Additional comparison on keypoint transfer in ShapeNet.
@@
Figure 24. More qualitative results of shape co-segmentation in ShapeNet.